OpenStudy (anonymous):

Algebra II Simplify square root 72 minus 3 times square root 12 plus square root of 192.

OpenStudy (anonymous):

factor out 12 under each radical. so sqrt(12*6)-3*sqrt(12)+sqrt(16*12). Factor out sqrt(12) to get sqrt(12)*(sqrt(6)+1)

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (anonymous):

If I factor out 12 from the square root of 72 I would get: 6 Square root of 2?

OpenStudy (anonymous):

\[6\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{72} = \sqrt{36*2}\] \[\Large \sqrt{72} = \sqrt{36}*\sqrt{2}\] \[\Large \sqrt{72} = 6\sqrt{2}\]

DivineSolar (divinesolar):

I mean if you simplify it down...\[6\sqrt{2} + 2\sqrt{3} \]

DivineSolar (divinesolar):

You can get your final answer out of this.

OpenStudy (anonymous):

Okay, as I solved the problem further I got that as my answer.

OpenStudy (anonymous):

Also, I need help with the following question: Simplify the expression fraction with numerator of the square root of negative four and denominator of the quantity three plus i minus the quantity two plus three times i.